For each of the following situations involving single amounts, solve for the unknown. Assume that interest is compounded annually. (i= interest rate, and n number of years) (EV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.) (Round your final answers to nearest whole dollar amount.) Present Value Future Value 58,000| 6.0% 2. $ 21,30272,000 18 11,718 $ 64,134 S 11,354 44,500| 10.0% 4.$ 64,134 $ 145,000 14 8.0%
Answer #1

Present Value

Future Value

i

n

1

$40,890

$58,000

6.0%

6

2

$21,302

$72,000

7.0%

18

3

$11,718

$44,500

10.0%

14

4

$64,134

$145,000

6.0%

14

5

$11,354

$19,458

8.0%

7

Explanation:

1.

Present Value = Future Value x PVIF (i, n)

Present Value = $ 58,000 x PVIF (6 %, 6)

                         = $ 58,000 x 0.7050 = $ 40,8902.

Present Value /Future Value = PVIF (i, n)

$ 21,302/$ 72,000 = PVIF (i, 18)

PVIF (i, 18) = 0.295861111 or 0.2959

PVIF (7%, 18) = 0.2959

3.

Present Value /Future Value = PVIF (i, n)

$ 11,718/$ 44,500 = PVIF (10 %, n)

PVIF (10 %, n) = 0.263325843 or 0.2633

PVIF (10 %, 14) = 0.2633

4.

Present Value /Future Value = PVIF (i, n)

$ 64,134/$ 145,000 = PVIF (i, 14)

PVIF (i, 14) = 0.442303448 or 0.4423

PVIF (6 %, 14) = 0.4423

5.

Future Value = Present Value x FVIF (i, n)

                       = $ 11,354 x FVIF (8 %, 7)

                       = $ 11,354 x 1.7138

                       = $ $19,458.49 or $ 19,458